Ajay has written a total of seven examinations in a sequence.He scored 90 marks in one exam & a total of 330 marks in the other six exams.If his average score in the first six exams is 60,find his score in the last exam.
A)35
B)30

Answer by richwmiller(17219)

90+330=t
420=t
6*60=360
420-360=60
He scored 60 on the last exam